In a bottling plant we need to test the pressure in abottle and the same have to filled in certian parmeter , they got two test,samling is done on 30 bottle after doing testin on first , the qulaity manager said the the test was not correct so he did the seocnd test on second machine ........... means on 30 first machine smaple is 181.5 and 181.7 how we will be testing hypothiese ?
it is derised that mean pressure strength is 175psi from the previous exeprience the stand deviation is more or less stable at 10.

Answer by stanbon(75887)

In a bottling plant we need to test the pressure in a bottle and the same have to filled in certian parameter , they got two tests sampling is done on 30 bottle after doing testing on first , the quality manager said the the test was not correct so he did the second test on second machine ...........
means on 30 first machine sample is 181.5 and 181.7 how we will be testing hypothises ?
it is desired that mean pressure strength is 175 psi from the previous experimemt; the standard deviation is more or less stable at 10.
---------------------
If I read you correctly:
1st test:
n = 30 bottles
mean = 181.5 with standard deviation of 10
-----------
2nd test:
n = 30 bottles
mean is 181.7 with standard deviation of 10
----
Desired pressure is 175.
===========================
Possible Hypothesis:
Ho: u = 175
H1: u > 175
-----
Assume that alpha = 5%
Using 1st test data the test statistic is:
t(181.5) = (181.5-175)/(10/sqrt(30)) = 3.5602
Then the p-value is P(t> 0.3.5602 when df=29) = 0.0006504
---
Conclusion based on the 1st test:
Since the p-value is less than 5%, reject Ho and conclude that
the pressure is statistically greater than 175.
-----------------------------------------------------
If you use the 2nd set of test results:
test stat = t(181.7) would give you even stronger
evidence that the pressure is greater than 175.
===================================================
Hope this helps.
Cheers,
Stan H.